OpenStudy (anonymous):

integrate x sin (2x^2) dx

sam (.sam.):

u-sub, u=2x^2 du=4xdx \[\text{}=\frac{1}{4}\int\limits \sin (u) \, du\] solve

sam (.sam.):

\[-\frac{\cos (u)}{4}+c\] \[-\frac{1}{4} \cos \left(2 x^2\right)+c\]

OpenStudy (anonymous):

integrate x sin (2x^2) dx =integrate sin (2x^2) d(1/2)x^2) =(1/2)(1/2)integrate sin(2x^2)d(2x^2) =-1/4 cos(2x^2) +C

OpenStudy (anonymous):

got it. thank you